OpenStudy (anonymous):

the expression 4.95+0.07x models a household's monthly long-distance charges, where x represents the number of minutes of long-distance calls during the month. what are the monthly charges for each number of long-distance minutes?? 10.) 73 minutes 11.) 29 minutes

OpenStudy (anonymous):

Are the 73 minutes and the 29 minutes multiple choice answers?

OpenStudy (anonymous):

no those are the two problems

OpenStudy (anonymous):

Do they want us to find the monthly charges for 73 minutes and 29 minutes??? Because that would make more sense

OpenStudy (anonymous):

what are the monthly charges for each number of long-distance minutes?

OpenStudy (anonymous):

sorry its taking me awhile to reply

OpenStudy (anonymous):

to answer your question yessss

OpenStudy (anonymous):

Okay, that's a lot better, I'm going to set up how to solve it and you solve it. I'll correct it.

OpenStudy (anonymous):

i got 10.) $366.46 and 11.) $145.58 for my answers

OpenStudy (anonymous):

Not quite, how did you come up with those answers?

OpenStudy (anonymous):

for number 10 I took 4.95+0.07=5.02(73)=$36.46 and for number 11. i took 4.95+0.07=5.02(29)=145.58 those numbers in my work i got from question and problem

OpenStudy (anonymous):

Where did you get 5.02 from?

OpenStudy (anonymous):

that's what the numbers add up and equal

OpenStudy (anonymous):

the numbers i added 5.02 is the result

OpenStudy (anonymous):

Oh, you added 4.95 and .07! They are not like terms because 0.07 is multiplied by "x" therefor you can't add them. Here try solving it this way: Plug in 73 for x. Making the equation: 4.95+0.07(73)

OpenStudy (anonymous):

they don't have o be like terms for 10 and 11

OpenStudy (anonymous):

oh i seee

OpenStudy (anonymous):

Yeah, and try 4.95+0.07(29) for the other one:)

OpenStudy (anonymous):

trying now. don't go i have more questions

OpenStudy (anonymous):

Wait, what were the answers you got?

OpenStudy (anonymous):

10.) 10.06

OpenStudy (anonymous):

Yes, don't forget that we are talking about money, so $10.06;)

OpenStudy (anonymous):

What about the next problem?

OpenStudy (anonymous):

$6.98

OpenStudy (anonymous):

Correct:)

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

No problem, ever need more help, contact me @JupitersBrain or become my fan lol. Goodluck!

OpenStudy (anonymous):

you leaving?

OpenStudy (anonymous):

Well, need more help?

OpenStudy (anonymous):

yes please??

OpenStudy (anonymous):

Okay, do the same thing you did. Create a new question and I'll view it.